# A Conceptual Framework For White Box Neural Networks

### v1

@@ Abstract
This paper introduces semantic features as a general conceptual framework for fully explainable neural network layers. A well-motivated proof of concept model for relevant subproblem of MNIST consists of 4 such layers with the total of ~4.8K learnable parameters. The model is easily interpretable, achieves human-level adversarial test accuracy with no form of adversarial training, requires little hyperparameter tuning and can be quickly trained on a single CPU. The general nature of the technique bears promise for a paradigm shift towards radically democratised and truly generalizable white box neural networks. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
Current state of the art (SOTA) neural networks are large, unwieldy black boxes offering little insight into their decision-making process, require prohibitive amounts of resources and are prone to adversarial attacks not seen during training [10, 3]. The latter implies that models rely on unintended features shared between train and test, giving false sense of generalisability [4, 6]. These issues drastically limit the usability of current generation of neural networks, especially in critical domains such as healthcare, cybersecurity or autonomous systems. Yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135k neurons. This discrepancy indicates a need for radically simplified AI.
This paper aims to lay foundation for a new paradigm for training models by introducing a conceptual framework for building white box neural networks. The emphasis is put on theoretical and qualitative aspects of proposed solutions and less on tweaking quantitative metrics. Following this spirit it’s important to train a proof of concept (PoC) network on Minimum Viable Dataset (MVD) - problem easy enough to get rid of all the obfuscating details but still hard enough to be relevant. MNIST dataset is falsely considered to be "solved" or "too easy" while SOTA models are hardly adequate in terms of their complexity, lack interpretability and are vulnerable to adversarial perturbations not seen during training [9]. This is true not only for the entire 10-digit dataset but even for certain pairs of digits [11] and therefore the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" (one of the most often confused pairs), which actually might be even harder than the entire MNIST in terms of average adversarial accuracy. This is not to say that similar results cannot be obtained on more complex datasets - preliminary experiments suggest that it is exactly the opposite - but in fact one of the goals of such selection of MVD is to make a clear and concise argument for the effectiveness and potential of presented method.
The structure of the paper is as follows:
@@
  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability and adversarial robustness.
@@ 2 Semantic features
Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [8]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias111by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 2.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel222can be made differentiable as in [7, 5].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices333preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images2
@@
    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls the non-zero element of $g$ by $p$ coordinates to the right, where $g$ is a single-entry vector of dimension $n$
@@ 2.2 Matching semantic features
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$. Therefore semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates444i.e. $f_{p}$ is a conjunction of lower-level features.. Therefore semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be "understood" by neural networks. This is especially apparent in the context of logical semantic features which allow to explicitly express logical sentences, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
  • $f=[1,0,-1,-1]$
  • $\mathbf{P}=\{0,1\}$
  • $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$
Semantic features capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time. As the remainder of the paper will show this turns out to be a sufficiently strong regularization - backpropagation through appropriately chosen $\mathbf{L}$ results in easily interpretable $f$ and $\mathbf{P}$.
@@
  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 555alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
@@ 3 Network architecture
The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $4.8$K parameters.
@@ 3.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH666the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 3.2 Convolutional Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches777a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
@@ 3.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small888close to the identity augmented matrix affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is located in the center of $28\times 28$ image, i.e. we zero-pad $f$ on every side before applying the affine transformations. Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
@@ 4.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was the Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability999mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
Reliability curveRefer to caption
Learning curveRefer to caption
@@ 4.3.1 Two Step Layer
Figure A show the initial and learned shapes of the Two Step function. Te function smoothly thresholds the input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
@@ 4.3.2 Convolutional Layer
Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned a meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
@@ 4.4 Decision boundary
Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and lables denote respectively:
@@
    top pose of affine features predicting label "3" (with the match value) [^10] These rows are actually overlaid with the middle row to make it easier to see what has been matched.
@@
    top pose of affine features predicting label "5" (with the match value)10
@@
5 Significance and further research
The importance of explainability is hard to overestimate. An automatic system discovering meaningful predictive patterns in data could significantly offload the task of finding the underlying causal structure of observed phenomena. This has a great potential to impact science, diagnostics, big data and many other fields. Combined with the demonstrated level of efficiency the resulting democratisation could open up exciting areas for growth worldwide, favouring diverse open source solutions and providing considerable level of robustness against centralisation.
The general nature of semantic features makes it reasonable to expect that extending these results to more complex datasets and domains should be relatively straightforward - but may require some honest engineering. The obvious ideas for further work include:
@@ 5 Significance and further research
  • design richer modes of spatial variation, e.g. hierarchical affine semantic features to capture variations of compound objects (objects consisting of affine parts are themselves affine if you "zoom out" appropriately);
  • find adequate semantic topology for color space;
  • integrate affine features with sliding window approach, i.e. implement weight sharing of similar features at distant locations;
  • define semantic features for text, e.g. to capture semantic invariance of permutations of words;
@@ 6 Acknowledgments
This research has been entirely self-funded - I saved money, left my job and took the risk of following independently a depressingly underfunded and underestimated line of research.
However it needs to be noted that I had the opportunity to do preliminary research on adversarial robustness as a half-time activity during my employment at MIM Solutions. Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract. It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
The story of this paper - should the results be recognized by the community - indicates a need for deep systemic and cultural changes that would empower individuals and smaller institutions to allow them to afford putting values and long-term informed risks over short-term profits. This is especially needed in larger public institutions like universities. Otherwise we might get stuck following "easy" directions which leads to a lethal combination of demoralisation and deterioration of quality.

### v2

# Towards White Box Deep Learning

@@ Abstract
This paper introduces semantic features as a candidate conceptual framework for building inherently interpretable neural networks. A proof of concept model for informative subproblem of MNIST consists of 4 such layers with the total of ~5K learnable parameters. The model is well-motivated, inherently interpretable, requires little hyperparameter tuning and achieves human-level adversarial test accuracy - with no form of adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
Machine learning systems used in safety-critical applications such as healthcare, military, autonomous vehicles or cybersecurity need to be reliable, i.e. they cannot fail unpredictably. The black box nature of state of the art (SOTA) deep neural networks (DNN) make it exceedingly hard to give such guarantees, which limits their usability to non-critical domains. The recent research in Explainable AI (XAI) focuses predominantly on post-hoc attribution methods which may provide explanations but, due to the complexity of SOTA, are hard to be converted to reliable fail-safes. This is a burning issue because of the wide susceptibility of DNNs to distribution shifts and adversarial attacks [18, 5] meaning that models can make arbitrarily incorrect predictions and their reported high test-time metrics don’t guarantee robust real-life performance111models often rely on spurious features that leak from test to train [6].. The recently popular Neurosymbolic AI [17] consists of higher-level methods that focus on augmenting neural network representations (machine perception) with background structured knowledge (reasoning) but don’t address the potential failures of the former.
Existing methods of improving DNN robustness are far from ideal. Adversarial Training [11] being one of the strongest empirical defenses against attacks [1] is slow, data-hungry [15, 8] and results in robust generalisation gap [13, 21] as models overfit to the train-time adversaries. Thus, the biggest challenge of empirical approaches is reliable evaluation of robustness as more and more sophisticated attacks break the previously successful defenses [3, 4].
To address these challenges the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [9]. However the exact certification is NP-Complete [20] and therefore has serious limitations in practice as it’s limited to relatively small problems and small networks and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
The highlighted issues are relevant for datasets as simple as MNIST [16] or even some of its binary subsets [19] - despite being "toy" and almost linearly separable problems, they still remain elusive in terms of robustness, rising further questions about the true generalisation capabilities of SOTA as no neural network model can be honestly considered reliable and adequate333in terms of internal complexity; SOTA models define inappropriately complicated functions. for this allegedly "simple" problem.
1.1 Proposed approach
This paper proposes an intermediate approach to DNN reliability and introduces a candidate theoretical framework for building white box neural networks. Building and analysing a white box DNN would be similar to writing and debugging a procedural program - it doesn’t have to come with theoretical certificates but needs to be easily understandable and internalizable by the programmer. A good software has to be designed well enough to account for reasonable failure scenarios, handle foreseeable exceptions and, should an unforseen bug happen, be amenable for fixes. If the design and inner mechanics of a DNN are sufficiently clear then working with such model wouldn’t be very different from maintaining a standard application.
Current SOTA models are large, unwieldy black boxes that require prohibitive amounts of resources; yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135K neurons! This scale seems easily maintainable with proper abstractions. After all, the core algorithms used today are mostly unchanged since ~30 years! What if there really exists a better way?
To find it we must start with the basics and take the agile approach. Focusing on the simplest relevant problem cuts off the obfuscating details and speeds up the experimentation phase. With that in mind the Minimum Viable Dataset (MVD) for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [19] and therefore might be considered harder than the entire MNIST in terms of average adversarial accuracy444the preliminary results on the full MNIST are discussed in Appendix C..
1.2 The core idea
This paper introduces semantic features as a general idea for sharing weights inside a neural network layer. The rigorous definition will be given in the next section. The concept is similar to that of "inverse rendering" in Capsule Networks [14] where features have many possible states and the best-matching state has to be found. Identifying different states by the subsequent layers gives rise to controlled dimensionality reduction. Thus semantic features aim to capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time. This is in fact a pretty strong regularization. As shown in this paper, choosing appropriate layers of semantic features for the MVD results in what can be considered as a white box neural network.
The remainder of the paper is structured as follows:
@@ 1.2 The core idea
  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and ablation;
  • Section 5. provides a short discussion of the results;
@@ 2 Semantic features
Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [14]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias555by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 2.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel666can be made differentiable as in [12, 7].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices777preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images6
@@
    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls elements of $g\in\mathbb{R}^{n}$ by $p$ coordinates to the right
@@ 2.2 Matching semantic features
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates888i.e. $f_{p}$ is a conjunction of lower-level features.. Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
  • $f=[1,0,-1,-1]$;           $\mathbf{P}=\{0,1\}$;           $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$.
@@
  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 999alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
@@ 3 Network architecture
The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters.
@@ 3.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH101010the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 3.2 Convolutional Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches111111a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
@@ 3.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small121212close to the identity augmented matrix affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image. Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
@@ 4.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability131313mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
Reliability curve
Learning curve
@@ 4.3.1 Two Step Layer
Figure A shows the initial and learned shapes of the Two Step function. The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
@@ 4.3.2 Convolutional Layer
Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
@@ 4.4 Decision boundary
Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and labels denote respectively:
@@
    top pose of affine features predicting label "3" (with the match value)141414these rows are actually overlaid with the middle row to make it easier to see what has been matched.
@@
    top pose of affine features predicting label "5" (with the match value)14
@@
4.5 Ablation study
Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%. The Convolutional Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
Interestingly the number of affine semantic features can be reduced from 8 to 2 with the cost of just ~2%pt adversarial accuracy. Logical Layer can be replaced by a simple linear probe, however it needs a proper initialization to avoid the risk of dead neurons (which is high because of the exceedingly small amount of neurons).
The Gaussian noise augmentation is vital for the model to experience pixel intensity variations during training (boosts adversarial robustness by ~62%pt).
Despite being primarily designed to withstand the $L_{\infty}$ attack the model has considerable robustness against the standard $L_{2}$ AutoPGD attack (eps=1.5, eps_step=0.1) with ~89% adversarial accuracy.
5 Discussion
We’ve built a white box neural network that is naturally robust, reliable and inherently interpretable. Thanks to the notion of semantic features the construction of the model was as straightforward as writing a fully explainable procedural program and so are the inner workings of the network.
Affine Layer is the core component of the model and the main reason for its interpretability. However if applied directly to the input it is easily fooled by attacks. It’s white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 6 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature). Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate. This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features. These layers are themselves compatible with the framework of semantic features which suggests a large untapped potential of the notion.
The Logical Layer has been developed as an attempt to formalize the idea of an abstract hidden layer sparse activation structure and in that regard it is the most inspiring; however it might turn out to be too generic (on MNIST the properly-initialized linear probe seems to be sufficient).
6 Further research
Future work would have to involve more complex datasets and more nuanced semantic features. For example notice the analogy between affine semantic features and the SIFT [10] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional) are essentially performing input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches parts-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
Other ideas for further work include:
@@ 6 Further research
  • study non-affine perturbations of features, e.g. projective perturbations;
  • find adequate semantic topology for color space;
  • define semantic features for text, e.g. capture semantic invariance of permutations of words;
@@ 7 Acknowledgments
This is work was done as an independent self-funded research.
However it needs to be noted that I had the opportunity to do preliminary investigation on adversarial robustness as a half-time activity during my employment at MIM Solutions. Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability of neural networks, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract. It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk and members of the reddit community for the great early feedback on the first version of this paper.